As studied, the average number of hours spent by senior high school students for their online classes a week is 25 hours with a standard deviation of 4 hours. Assuming that the study is is true and the data is normally distributed.

1
Expert's answer
2022-02-28T18:07:19-0500

Let "X=" the number of hours spent by senior high school students for their online classes a week: "X\sim N(\mu, \sigma^2)"

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=25\ h, \sigma=4\ h."

a.


"P(26<X<29)=P(X<29)-P(X\le26)"




"=P(Z<\dfrac{29-25}{4})-P(Z\leq\dfrac{26-25}{4})"




"=P(Z<1)-P(Z\leq0.25)"




"\approx0.8413447-0.5987063\approx0.2426"



b.


"P((X<21)\cup(X>30))"




"=P(X<21)+1-P(X\leq30)"




"=1+P(Z<\dfrac{21-25}{4})-P(Z\leq\dfrac{30-25}{4})"




"=1+P(Z<-1)-P(Z\leq1.25)"




"\approx1+0.158655-0.894350\approx0.2643"



с.


"\bar{X}\sim N(\mu, \sigma^2/n)"

Given "n=12"


"P(\bar{X}>24)=1-P(\bar{X}\le 24)"




"=1-P(Z\le\dfrac{24-25}{4/\sqrt{12}})"




"\approx1-P(Z\le-0.866025)"




"\approx0.8068"